factor the left side
x^2+4x+4=7

Found 2 solutions by stanbon, venugopalramana:
Answer by stanbon(75887)

factor the left side
x^2+4x+4=7
(x+2)^2=7
Take the sqrt of both sides to get:
x+2 = sqrt7 or x+2=-sqrt7
x=-2+sqrt7 or x=-2-sqrt7
Cheers,
Stan H.

Answer by venugopalramana(3286)

factor the left side
x^2+4x+4=7
LHS=(X)^2+2*(X)(2)+(2)^2=(X+2)^2....(A+B)^2=A^2+2*A*B+B^2